1. Exercise 17.1

A firm has the opportunity to invest in a project having an initial outlay of $20,000. Net cash inflows (before depreciation and taxes) are expected to be $5,000 per year for five years. The firm uses the straight-line depreciation method with a zero salvage value and has a (marginal) income tax rate of 40 percent. The firm’s cost of capital is 12 percent.

a) What is the internal rate of return (IRR) for the project? ( Answer-------------)(Hint: Present value of an ordinary annuity of $1 for 5 years is $4.3295 at the discount rate of 5% and $4.4518 at the discount rate of 4%.)

b) What is the net present value (NPV) of cash flows from the project? (answer----------)

c) Should the firm accept the project?

Yes

No
Answer #1

Answer:

Given

Project Initial cost I=$20000

Project life N= 5

Depreciation each year D=20000/5=$4000

Before tax Net cash inflow C=$5000

Taxable income =C-D=5000-4000=1000

Tax rate =40%

tax =40%*1000=$400

So after tax cash flow P= 5000-400=$4600

1) Let r be the cost of capital so at IRR

I=P*(1-(1+r)^-N)/r

20000=4600*(1-(1+r)^-5)/r

So solving for r we get

r=4.85%

2) Given cost of capital r=12%

NPV=P*(1-(1+r)^-N)/r-I

NPV=4600*(1-(1+12%)^-5)/12%-20000=-$3418

3)

Since we are getting negative NPV so we don't accept this project.